OpenStudy (anonymous):

An acute angle θ is in a right triangle with cos θ = nine tenths. What is the value of sec θ?

OpenStudy (lochana):

can you write sec(theta) in terms of cos(theta)?

OpenStudy (lochana):

well, there is a co-relation between sec and cos. \[\sec \theta = \frac{ 1 }{ \cos }\] now you should be able find sec theta

OpenStudy (lochana):

do you feel that you can do it yourself.

OpenStudy (lochana):

@zoeg are you following me:)